OpenStudy (anonymous):

Good Morning Y'all Please help, a random variable r is normally distributed with a mean of 7 and a standard deviation of 2. Find the value of W, so that P(W

OpenStudy (anonymous):

Howdy, Cola =)

OpenStudy (anonymous):

hey hey, I hope you can help me with this

OpenStudy (anonymous):

So, first of all, a Z-score of a number is pretty simply just describing how many Standard Deviations from the mean that number is. What is the z-score for 10?

OpenStudy (anonymous):

Since the mean is 7 and the SD is 2.

OpenStudy (anonymous):

ill let you in on how I got my magical ( probly wrong) anwer NormalCDF(-1*10^99,10,7,2) = .9332 .9332 - .8913 = 0.0418 invNorm(0.0418,7,2) = 3.54

OpenStudy (anonymous):

I cant just find a z score because it is not on a solid "line" 7 + 2 = 9

OpenStudy (anonymous):

Okay, well your method is a bit strange to me, haha, but I've looked over it, and it seems to be valid.

OpenStudy (anonymous):

Thank you so very much so my answer is not my r but rather my w right ???

OpenStudy (anonymous):

Well, I'm not sure what your complaint is. "I cant just find a z score because it is not on a solid "line" 7 + 2 = 9"

OpenStudy (anonymous):

Let me try my own method real quick and see what I get.

OpenStudy (anonymous):

 see there is no way to get a z from just the drawing and the given so we need a calculator :)